Find the sum of n terms of the series whose r term is a²r
 

Grade:12

2 Answers

Arun
4 years ago
Dear Kavya
 
sum of n terms = a^2 ∑r = a^2 (n(n+1)/2)
 
In case of any difficulty, Please feel free to ask again.
Vikas TU
4 years ago
Hiii 
First of all clearify , It is AP or GP , 
Wthout this , we can not solve this question , 
Feel free to ask .
Good luck 
Cheers